OpenStudy (anonymous):

Can anyone help find the inverse of the fxn: y=x^(2)+4x-6

OpenStudy (anonymous):

it is not a one to one function, so it is unlikely that it has an inverse

OpenStudy (anonymous):

there are 2 inverses, depending on the chosen domain

OpenStudy (anonymous):

I have the answer, I just don't understand how to get there

OpenStudy (anonymous):

ok then we can solve via completing the square \[y+6=x^2+4x\]\[y+6+4=(x+2)^2\] \[y+10=(x+2)^2\] then \[x+2=\pm\sqrt{y+10}\] and so \[x=-2\pm\sqrt{y+10}\]

OpenStudy (anonymous):

where did the +4 come from?

OpenStudy (anonymous):

because \(x^2+4x\neq (x+2)^2\) however \((x+2)^2=x^2+4x+4\) so when we replace \(x^2+4x\) by \((x+2)^2\) we were adding 4 therefore you have to add 4 to the other side as well

OpenStudy (anonymous):

aka "completing the square"

OpenStudy (anonymous):

ahh, thank you, I never would have thought to complete the square

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

satellite73 do you know the domains?

OpenStudy (anonymous):

sure it is given by the plus minus part

OpenStudy (anonymous):

the question is worded strangely, it asks which domains lead to the two inverses

OpenStudy (anonymous):

vertex of the parabola \(y=x^2+4x-6\) is \((-2,-10)\) so if \(x<-2\) the "function" is decreasing, while if \(x>-2\) it is increasing

OpenStudy (anonymous):

Ok, thank you

OpenStudy (anonymous):

therefore if \(x<-2\) the inverse is the one with the negative radical, if \(x>-2\) use the one with the positive radical yw